OpenStudy (anonymous):

If y varies directly as x, and y is 20 when x is 4, what is the constant of variation for this relation?

OpenStudy (anonymous):

5 heres where https://answers.yahoo.com/question/index?qid=20150318104659AAxf81F

OpenStudy (haseeb96):

y directly as x y= k x where k is constant now y=20 and x=4 put it in the above equation 20= k (4) k=20/4 k=5 this is the correct answer

OpenStudy (haseeb96):

@maxz18 did u get my point?